The demand for schedule for beer is Xd=25-P, where Xd is the quantity demanded of beer in millions of barrels per year and P is price in dollars per barrel.




If the supply curve for beer is Xs = -20 +4P, what is the equilibrium price of a barrel of beer?




Calculate and interpret price elasticity of demand and supply at the equilibrium point.




What would the effect on the price of a barrel of beer if a tax of $ 4 per barrel is imposed by the government?

1
Expert's answer
2022-02-20T16:27:36-0500

At equilibrium, "Xd= Xs"

"25-p=4p-20"

"4p=45"

"P=9"

"Q=25+P=25-9=16"

Price elasticity of demand at equilibrium"=\frac{price at equilibrium}{Quantity at Equilibrium}×\frac{dQ}{dP}"

"\frac{dQ}{dP}=25-1=24"

Therefore;

"P.E.D=\frac{9}{16}×24= 13.5"

Beer is an elastic product since it P.E.D is greater than 1.

"Change InQuantity= P.E.D\u00d7\u2206In Price."

"=4\u00d713.5=54"

Thus the quantity supplied will change by 54.